How much charge does a 16 V battery have to supply to fully charge a 3.0 μF capacitor and a 5.2 μF capacitor when they’re in parallel? =130 a. How much energy does the battery have to supply in this case?
Answer #1

capacitors are in parallel, so

equivalent capacitance = 5.2 uF + 3.0 uF = 8.2 uF

charge Q = C*V = 8.2*10^(-6)*16 = 13.12*10^(-5) C

energy = 0.5*CV^2 = 0.5*8.2*10^(-6)*16^2 = 0.00105 J = 1.05*10^(-3) J